Part A stomach acid, 1.1×10−13M Express your answer to two significant figures and include the appropriate...

Part A

stomach acid, 1.1×10−13M

Express your answer to two significant figures and include the appropriate units.

Part B

Part complete

urine, 8.7×10−9M

Express your answer to two significant figures and include the appropriate units.

Part C

Part complete

orange juice, 4.3×10−11M

Express your answer to two significant figures and include the appropriate units.

Part D

Part complete

bile, 1.8×10−6M

Express your answer to two significant figures and include the appropriate units.

Please explain how to enter these in calculator.
Answer #1

Answer - (a] Given [on) = 1.1 хон Ги, о*] ? We know that, рон - - Cog Lon] — Also, РИ + pH - Ч Put ① in ② ри = 1ч+ оз Con) (3

1. 13) (14 -- los Put - the value, [H3ot] =10 10 ( 14 - 12 96) [Hoot] = 1 [Hot] = 9.18/02M) Part B. Given [on] = 8.7810-9 M [